OpenStudy (anonymous):

The length of a diagonal of a square is 8.5 inches. Find the area of the square.

terenzreignz (terenzreignz):

Were you given a formula as to how to get the area of the square given the length of its diagonal?

OpenStudy (anonymous):

i was never given such a formula, but it cannot be that hard to figure out if the side of a square is \(a\) then the area is \(a^2\) and by pythagoras the diagonal is \(a\sqrt{2}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

this, since \(a^2+a^2=h^2\) so \(2a^2=h^2\) making \(h=a\sqrt{2}\)

OpenStudy (anonymous):

how to i get an \(a^2\) from \(a\sqrt{2}\)? square it, and divide by \(2\)

OpenStudy (anonymous):

so you job is to take \(8.5\) and square it, then divide the result by \(2\)